# GDSG: Graph Diffusion-based Solution Generation for Optimization Problems in MEC Networks

### v1

@@ V-A1 Datasets
We tentatively tried using GEKKO [48] solver to generate non-graph data for a single server, but the solution time for a single solve of a four-node MINLP problem reached several seconds. Moreover, GEKKO cannot handle graph structure data. As a result, we use a heuristic algorithm with polynomial time complexity to generate suboptimal datasets for the computation offloading problem in Sec. II. The Minimum Cost Maximum Flow (MCMF) algorithm, applied to random heuristic initialization of edge weights, efficiently finds optimal edge selection results and provides high-quality suboptimal solutions (solving a single instance with $80$ nodes in $0.6$sec). Since an optimal solver is not available, we use exhaustive methods to generate small-scale ground-truth datasets but the time complexity is very high (over $30$sec for a single instance of $80$ nodes).
@@ V-A3 Metrics
V-A4 Baselines
We use five baseline methods: the heuristic algorithm with the MCMF algorithm, which is the tool used to generate the suboptimal datasets for GDSG and other baselines, and the corresponding code has also been open-sourced; the discriminative GNN (DiGNN) with the same neural network structure as our GDSG; and MTFNN [22], an off-the-shelf multi-task learning DNN method; DIFUSCO [36] and T2T [37] are also included as latest known methods that use diffusion generation for combinatorial optimization. For all experiments, each hyperparameter for a given model and training set is fixed by default (details are in the open source). For the following sections, the neural network settings for GDSG and DiGNN are identical. GDSG shares the same diffusion-related hyperparameter settings as DIFUSCO and T2T. All dataset generation is done on an Intel i7-13700F, and all model training and testing are performed on an RTX 4090.
@@ V-A4 Baselines
V-B Empirical Verification for Theory
Based on (Eq. (12), Eq. (13)), we demonstrate $\mathbb{E}(q(\theta^{n}))$ for three cases in both discrete and continuous diffusion: $N=1,10,20$. We set the possible values for the unknown variables in the equation under the assumptions in Sec. III-B. For discrete solutions, we set $\epsilon\!=\!0.1$. For continuous solutions, we uniformly set $y_{i}=1$, $\epsilon\!=\!0.1$, $\gamma_{i}\!=\!0.04$, and $\delta=0.1$. The expected value of $\mathbb{E}(q(\theta^{n}))$ is then computed, resulting in Fig. LABEL:fig_empirical_ver. We use dotted lines of the same color as the curves to mark the corresponding lower bound on the number of samples $n$ in Eq. (12) and Eq. (13). Here we consider $\mathbb{E}(q(\theta^{n}))>0.95$ as $\mathbb{E}(q(\theta^{n}))\approx 1$. As $n$ increases, the expectation approaches $1$. Although some unknown variables are assigned assumed values, Fig. LABEL:fig_empirical_ver visually illustrates the convergence process of sampling from the high-quality solution distribution and the associated influencing factors.
Additionally, for the unknown variables in Eq. 9, which refer to the model’s generation data variance, we use empirical calculations based on the existing model outputs. The trained model is used to repeat 1000 inferences on the same input and count the variance of the output $\theta$s, which is on the order of $10^{-4}$. Therefore, $\frac{\sigma^{2}}{\epsilon^{2}}$ can be small, e.g., $\frac{\sigma^{2}}{\epsilon^{2}}=0.01\ if\ \sigma^{2}=0.0001,\epsilon=0.1$. Although this is not a perfect theoretical conclusion, it demonstrates the small, even almost negligible, gap between the learned distribution and the optimal distribution.
@@ V-B Empirical Verification for Theory
cells = c, hlines, vlines, Evaluation Dataset & MTFNN HEU DiGNN DIFUSCO T2T GDSG (Ours)
@@
V-C Multi-Task Orthogonality
As we try to address discrete diffusion and continuous diffusion in a single model, we should ensure the gradient vectors of the two tasks are orthogonal so there is no negative competition between the two tasks during training [57, 58]. The non-orthogonality of the gradient vectors of two tasks often leads to inferior convergence. Considering this kind of risk, we track the gradient directions of the related parameters for both tasks throughout the training process of our model and the baseline, DiGNN, with the same neural network structure. Specifically, we perform backpropagation of the loss for each task at each training iteration across multiple epochs. Then we extract a high-dimension vector, which contains the information from the parameter updating gradient, of each parameter of representative network layers (input layer, edge feature embedding layer, node feature embedding layer, etc). Then we compute the cosine of the angle between the update vectors of the two tasks to quantify their orthogonality. The results are shown in Fig. LABEL:fig_gradient_ana.
It is observed that in GDSG, the two tasks show a high degree of orthogonality with cosine values approaching $1$ and orthogonal gradient proportions approaching $100\%$, which is in sharp contrast to the baseline. For DiGNN, it shows poor orthogonality with cosine values far away from 1 and orthogonal gradient proportions always less than $10\%$. Furthermore, GDSG has a small variance (the maximum magnitude is $4$) while DiGNN has a large variance (the minimum magnitude is $7$), which means that the gradient angles of GDSG are more concentrated in orthogonality. This advantage is benefit from the configuration of the loss function of diffusion. Therefore, the simultaneous training of the two tasks in a single GDSG model does not impede each other, which avoids any degradation of model performance.
V-D GDSG vs DiGNN on Effectiveness and Generalization
[…]
@@ V-E Comparison with Baseline Schemes
DIFUSCO performs relatively worse than GDSG, with the performance gap primarily due to the lack of consideration for padding edges in the GNN. On the other hand, T2T introduces automatic gradient computation of the target optimization function’s gradient during the denoising process. While this is computable for TSP and MIS problems in the original work, our computation offloading objective function is non-differentiable, making gradients unavailable. As a result, T2T is not applicable to MEC network optimization problems like ours.
VI Conclusion and discussion
In this study, we introduce GDSG, a model for generating solutions based on graph diffusion, especially delivering optimization solution generation capabilities that are robust to training data. We focus on NP-hard optimization problems in MEC networks that lack effecient approximation algorithms, using the widely prevalent multi-server multi-user computation offloading problem as a concrete scenario. Our model converts network optimization problems into distribution learning through a defined parameterization of the solution space, which supports learning from suboptimal training sets that can be obtained efficiently. The paper provides important findings and explanations that advocate the application of generative models in solution generation. This application offers a cohesive problem formulation and model framework for network optimization. Furthermore, we construct a database for the multi-server multi-user computation offloading (MSCO) optimization problem, including suboptimal training sets, optimal training sets for a small number of small-scale problems, and optimal test sets. Through our experiments, GDSG has shown superior performance compared to baseline methods, whether on the suboptimal training set or the optimal training set. The proposed GDSG approach disrupts the current trend in learning-based network optimization methods that rely heavily on extensive ground-truth datasets, thereby reducing the reliance on real optimal datasets.
We have considered some future directions of this work, one of which is ensuring that generated solutions satisfy the constraints of optimization problems through more explicit mechanisms. Additionally, the presence of numerous unknown variables in the theoretical framework means that comprehensive conclusions on convergence or data requirements remain room for further investigation.

### v3

@@ V-A1 Datasets
We tentatively tried using existing numerical solvers such as Gurobi and GEKKO to generate non-graph data for a single server, but the solution time for a single solve of a four-node MINLP problem reached several seconds. Moreover, it is difficult for Gurobi and GEKKO to handle graph structure data. As a result, we use a heuristic algorithm with polynomial time complexity to generate suboptimal datasets for the computation offloading problem in Sec. II. The minimum cost maximum flow (MCMF) algorithm, applied to random heuristic initialization of edge weights, efficiently finds optimal edge selection results and provides high-quality suboptimal solutions (solving a single instance with $80$ nodes in $0.6$sec).
Specifically, MSCO can be formulated as a graph-level mixed-integer programming problem, where the discrete offloading decisions represent the data flow from users to servers. This formulation allows the problem to be solved using the MCMF algorithm. Initially, a heuristic is used to allocate computational resources and establish the continuous solution, i.e., $A_{i}=\frac{V_{i}}{\sum_{l,\Upsilon(E_{l},U_{j})\Omega(E_{l},S_{k})=1}V_{j}},\forall i\in\{1,\dots,M\},\ k\in\{1,\dots,K\}$. By applying additional randomness, we can compute the costs for both local execution and offloading, which allows us to determine the outbound flow costs for users. To meet the integer cost requirement of the MCMF algorithm, we retain three decimal places in the real-valued costs and scale these values by $1000$. The MCMF algorithm then computes the minimum-cost flow while ensuring maximum flow throughput. Our topology design guarantees that the maximum flow in the MCMF formulation always matches the number of computational tasks. This effectively reduces the problem to a straightforward minimum-cost flow, optimizing the offloading decision with minimal computation cost. By repeatedly initializing the computational resource allocation randomly, solving the MCMF, and selecting the best solution, we efficiently generate high-quality suboptimal solution samples. Since an optimal solver is not available, we use exhaustive methods to generate small-scale ground-truth datasets but the time complexity is very high (over $30$sec for a single instance of $80$ nodes).
@@ V-A3 Metrics
V-A4 Baselines
We use five baseline methods: the heuristic algorithm with the MCMF algorithm (HEU), which is the tool used to generate the suboptimal datasets for GDSG and other baselines, and the corresponding code has also been open-sourced; the discriminative GNN (DiGNN) with the same neural network structure as our GDSG; and MTFNN [38], an off-the-shelf multi-task learning DNN method; DIFUSCO [43] and T2T [44] are also included as latest known methods that use diffusion generation for combinatorial optimization. For all experiments, each hyperparameter for a given model and training set is fixed by default (details are in the open source). In the following sections, the neural network settings for GDSG and DiGNN are identical. GDSG shares the same diffusion-related hyperparameter settings as DIFUSCO and T2T. All dataset generation is done on an Intel i7-13700F, and all model training and testing are performed on an RTX 4090.
V-B Empirical Verification for Theory
This experiment aims to clarify the theoretical concepts discussed in Section III-B by demonstrating the expected probability of achieving an optimal solution through repeated sampling with specified hypothetical parameters. It also aims to illustrate the feasibility of learning the solution distribution necessary for reaching the optimal solution and analyzing key factors that influence the expected hit rate. These factors include the quality of suboptimal solutions within the training set and the dimensionality of the solution variables.
Based on (Eq. (12), Eq. (13)), we demonstrate $\mathbb{E}(q(\boldsymbol{\theta}^{n}))$ for three cases in both discrete and continuous diffusion: $N=1,10,20$. We set the possible values for the unknown variables in Eq. ((9), Eq. (10)) under the assumptions in Sec. III-B. For discrete solutions, we set $\boldsymbol{\epsilon}\!=\!0.1$. For continuous solutions, we uniformly set $y_{i}=1$, $\boldsymbol{\epsilon}\!=\!0.1$, $\boldsymbol{\gamma}_{i}\!=\!0.04$, and $\delta=0.1$. The expected value of $\mathbb{E}(q(\boldsymbol{\theta}^{n}))$ is then computed, resulting in Fig. LABEL:fig_empirical_ver. We use dotted lines of the same color as the curves to mark the corresponding lower bound on the number of samples $n$ in Eq. (12) and Eq. (13). Here, we consider $\mathbb{E}(q(\boldsymbol{\theta}^{n}))>0.95$ as $\mathbb{E}(q(\boldsymbol{\theta}^{n}))\approx 1$. As $n$ increases, the expectation approaches $1$. Although some unknown variables are assigned assumed values, Fig. LABEL:fig_empirical_ver illustrates the convergence process of sampling from the high-quality solution distribution and the associated influencing factors.
Additionally, for the unknown variables in Eq. (9), which refer to the model’s generation data variance, we use empirical calculations based on the existing model outputs. The trained model is used to repeat 1000 inferences on the same input and count the variance of the output $\boldsymbol{\theta}$s, which is on the order of $10^{-4}$. Therefore, $\frac{\boldsymbol{\sigma}^{2}}{\boldsymbol{\epsilon}^{2}}$ can be small, e.g., $\frac{\boldsymbol{\sigma}^{2}}{\boldsymbol{\epsilon}^{2}}=0.01\ if\ \boldsymbol{\sigma}^{2}=0.0001\ \boldsymbol{\epsilon}=0.1$. Although this is not a perfect theoretical conclusion, it demonstrates a small, even almost negligible, gap between the learned distribution and the optimal distribution.
V-C Multi-Task Orthogonality
As we address discrete diffusion and continuous diffusion in a single model, we should ensure the gradient vectors of the two tasks are orthogonal so there is no negative competition between the two tasks during training [67, 68]. The non-orthogonality of the gradient vectors of two tasks often leads to inferior convergence. Considering this challenge, we track the gradient directions of the related parameters for both tasks throughout the training process of our model and the baseline, DiGNN, with the same neural network structure. Specifically, we perform backpropagation of the loss for each task at each training iteration across multiple epochs. Then, we extract a high-dimensional vector that captures the information from the parameter-updating gradient for each parameter of the representative network layers. This includes the node feature embedding layer, the edge feature embedding layer, the three edge convolutional layers, and the learnable output normalization layer of the last graph convolutional module. Finally, we compute the cosine of the angle between the update vectors of the two tasks to quantify their orthogonality. The results are shown in Fig. LABEL:fig_gradient_ana.
It is observed that in GDSG, the two tasks show a high degree of orthogonality with cosine values approaching $1$ and orthogonal gradient proportions approaching $100\%$, which is in sharp contrast to the baseline. For DiGNN, it shows poor orthogonality with cosine values far away from 1 and orthogonal gradient proportions always less than $10\%$. Furthermore, GDSG has a small variance (the maximum magnitude is $4$) while DiGNN has a large variance (the minimum magnitude is $7$), which means that the gradient angles of GDSG are more concentrated in orthogonality. This advantage is benefit from the configuration of the loss function of diffusion. Therefore, the simultaneous training of the two tasks in a single GDSG model does not adversely affect each other, which avoids any degradation of model performance.
@@ V-C Multi-Task Orthogonality
V-D GDSG versus DiGNN on Effectiveness and Generalization
[…]
@@ V-D GDSG versus DiGNN on Effectiveness and Generalization
colspec=X[0.85,l] X[0.5,c] X[0.75,c] X[0.45,c] X[0.45,c] X[0.55,c] X[0.4,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, Evaluation Dataset & MTFNN HEU                       (dataset generator) DiGNN DIFUSCO T2T                    (not applicable) GDSG (Ours)
@@ V-E Comparison with Baseline Schemes
DIFUSCO performs relatively worse than GDSG, with the performance gap primarily due to the lack of consideration for padding edges in the GNN. Moreover, T2T introduces automatic gradient computation of the target optimization function’s gradient during the denoising process. While this is computable for TSP and MIS problems in the original work, our computation offloading objective function is non-differentiable, making gradients unavailable. As a result, T2T is not applicable to MEC optimization problems.
V-F Complexity Analysis
To facilitate the application of the GDSG model, we present a complexity analysis. Our model primarily features a graph convolutional neural network with a gating mechanism [65]. It incorporates input layers for edge feature embeddings, node feature embeddings, and time embeddings, along with output layers designed for both continuous and discrete solutions.
Let $V$ and $E$ denote the number of nodes and edges in the input graph, respectively. Let $S$ as the number of graph convolutional modules, $h$ as the hidden vector dimension, $d_{v}$ as the input node feature dimension, and $d_{e}$ as the input edge feature dimension. The time complexity and space complexity of the temporal embedding input layer are $O(h+\frac{3}{4}h^{2})$ and $O(2h)$, respectively. For the node feature embedding layer, the time and space complexity are $O(d_{v}h+h^{2})$ and $O(d_{v}+h)$. Similarly, the time and space complexity of the edge feature embedding layer are $O(d_{e}h+h^{2})$ and $O(d_{e}+h)$. For a single graph convolutional module, the time complexity and space complexity are $O(8h^{2}+2V+E)$ and $O(6h)$, respectively. During the graph convolution process, the time embedding layer has a time complexity of $O(\frac{1}{2}h^{2})$ and a space complexity of $O(\frac{1}{2}h)$, while each module’s output layer has a time complexity of $O(E+h^{2})$ and a space complexity of $O(E+h)$.
[…]
VI Conclusion
In this study, we have introduced GDSG, a model designed for generating solutions based on graph diffusion, particularly focused on solving optimizing problems that remain robust to the limitations of training data. Our primary emphasis is on NP-hard optimization problems in MEC networks that currently lack efficient approximation algorithms, with the multi-server multi-user computation offloading (MSCO) problem serving as a specific example. The proposed GDSG model transforms network optimization problems into a framework of distribution learning by systematically parameterizing the solution space. This allows us to efficiently learn from suboptimal training sets, which can be more easily gathered compared to optimal ones. Our findings highlight the potential of generative models in solution generation, offering a unified structured model framework for solving network optimization problems. The results of GDSG highlight a pathway that avoids the high complexity constraints of existing numerical algorithms while significantly enhancing the robustness of training data and the application efficiency of AI-based methods.
Additionally, we have developed a comprehensive database for the MSCO optimization problem. This includes suboptimal training sets, a limited number of optimal training sets for smaller-scale problems, and optimal testing sets. With the current experimental results, we believe that cross-scale generalization and model simplification through distillation could be two promising directions for future research.